# Distributed ADMM Approach for the Power Distribution Network Reconfiguration

### v1

@@ I Introduction
PDNR methods can be categorized into heuristic [2, 3, 4, 5], meta-heuristic [6, 7, 8] and mathematical optimization-based approaches [9, 10, 11]. Various methods were employed to address the PDNR problem by utilizing centralized methods [12]. These methods gather all network data and transmit it back to the control center to determine the reconfiguration solution. However, this approach is vulnerable to single-point failures and cyber-attacks. Furthermore, it can be computationally expensive and unsuitable for large-scale systems. An alternative approach is to use distributed (or decentralized) algorithms. In distributed algorithms, computing agents are only required to share limited amounts of information with a specific subset of other agents. This can enhance cybersecurity and reduce the costs associated with the necessary communication infrastructure. Additionally, distributed algorithms demonstrate robustness in the face of individual agent failures and enable the integration of renewable energy into the grid. Moreover, they can execute parallel computations, which positions them as potentially superior to centralized algorithms in terms of solution speed and the maximum problem size they can address [13].
@@
In the context of the distributed approach for the PDNR problem, the multi-agent approach has been suggested in various studies, including [14, 15, 16]. However, these approaches often require agents, or some of them, to have access to all information, which presents challenges related to data privacy and vulnerability to single-point failures. Moreover, they typically overlook the development of scalable solutions for large-scale distribution networks. Another approach to address the distributed PDNR problem is the Alternating Direction Method of Multipliers (ADMM) used in [17, 18, 19].
@@
The ADMM algorithm is a primal-dual splitting method initially designed to solve convex optimization problems [20, 21]. ADMM as a general heuristic for solving non-convex problems is mentioned in [20, Ch 9] and this idea is further explored in [22, 23]. Recently, it has garnered attention as a method for discovering approximate solutions for NP-hard problems. Utilizing ADMM for mixed-integer problems offers a significant benefit by requiring only a method to compute the projection onto the discrete constraint set, either exactly or approximately if the computation is costly [22]. While there is no guarantee that ADMM-based methods converge to a global solution for non-convex problems, they are still useful in finding high-quality local solutions.
ADMM algorithm has proven to be a powerful tool for solving the distributed optimal power flow problem by decomposing it into smaller, more manageable subproblems. Studies such as [24, 25, 25] demonstrate ADMM’s effectiveness in improving computational efficiency and system resilience without relying on centralized control. Recent works, including [26, 27] further highlight ADMM’s role in enhancing the robustness and scalability of power system optimization in distributed and cyber-physical contexts.
@@
Regarding the application of ADMM in PDNR problems, a distributed ADMM-based algorithm was presented in [17]. However, it can lead to infeasible solutions due to the lack of a guaranteed radial network structure, as only projections onto the binary set were considered. To address this issue, an enhanced ADMM-based algorithm was proposed in [18]. This algorithm incorporates a proximal operator with a penalty parameter based on residual errors to handle binary variables. Nevertheless, there is still no guarantee of achieving a radial solution since the projection step remains on the binary set. In [19], additional conditions were introduced in the projection step to ensure radiality. The authors combined the algorithms from [17] and [18] with their own, selecting the optimal (feasible) solution among the three. Notably, the examination of radiality constraints on binary variables occurs only after the algorithm’s convergence.
The Minimum Spanning Tree (MST) problem is a fundamental challenge in combinatorial optimization, requiring the identification of a tree that spans all vertices in a connected undirected graph while minimizing the total weight of its edges. In the context of directed graphs, the analogous concept to a tree is an arborescence [28], rooted at a specific node, ensuring a directed path from the root to any other node. This problem is known as the Minimum Weight Rooted Arborescence Problem (MWRAP) [28].
Various heuristic algorithms address the PDNR problem by utilizing solutions to the MST problem [2, 29, 30, 31, 32, 33]. However, these algorithms are centralized and rely on specific assumptions grounded in engineering knowledge of the distribution network. These assumptions include the anticipation that losses are near-optimal when all tie-switches are closed, forming a weakly meshed network. Additionally, they assume a moderate balance in load distribution among feeders and a voltage profile close to the best possible [29].
Contributions: This work makes the following contributions:
  • We present a distributed ADMM-based algorithm that improves upon the methods introduced in [17, 18, 19], where relaxation techniques were employed. Unlike these previous approaches, which relax the binary variable, our method utilizes a natural variable substitution. Numerical simulations demonstrate the superiority of our approach compared to the relaxation techniques.
  • We demonstrate that the projection step onto the set of radiality constraints in the ADMM algorithm is equivalent to solving an MWRAP. The specific weights in this context are determined by the flow, which is the solution to a linearly constrained convex quadratic programming problem, and the update of the ascent dual variable. This approach differs from previous works [17, 18, 19], which only project onto the binary set. Such binary projections may result in infeasible solutions.
@@ II The Optimization Problem for the Simplified DistFlow Model
Bold symbols consistently represent vectors or matrices. If $\boldsymbol{x},\boldsymbol{y},\boldsymbol{z}\in\mathbb{R}^{n}$ then $\boldsymbol{z}\in[\boldsymbol{x},\boldsymbol{y}]$ means that $z_{i}\in\left[x_{i},y_{i}\right]$ for all $i\in\{1,\ldots,n\}$. Moreover, $[\boldsymbol{x}]_{e}$ denotes the $e$-th element of the vector $\boldsymbol{x}$. The vectors $\boldsymbol{0}_{n}$ and $\boldsymbol{1}_{n}$ denote vectors of size $n$ consisting entirely of zeros and ones, respectively.
@@
$\mathbb{S}=\left\{\boldsymbol{b}\in\{0,1\}^{|\mathcal{A}|}:\ \boldsymbol{b}\ \text{forms an arborescence rooted at }g\right\},$
@@
$[\boldsymbol{A}\boldsymbol{U}]_{e}=U_{i}-U_{j},\quad\text{if }e=(i,j)\in\mathcal{A}.$
Note that a feasible solution $\boldsymbol{b}$ to the above problem defines an arborescence $G_{\boldsymbol{b}}\subseteq G$ (a directed tree) rooted at $g$, whose arcs are the closed switches. The reader can refer to [28, Subsection 2.2] for a detailed definition of an arborescence. Throughout this paper, we assume the feasibility of problem (6), meaning that at least one optimal solution exists.
@@
The voltage drop constraint (3) has been multiplied component-wise with $\boldsymbol{b}$ in (6), making it active solely when $b_{ij}=1$. This is different from the “big-$M$” method presented in [10, 36] which includes or excludes these constraints using the inequality:
@@ III-A Review of the MWRAP
Given a directed graph $G=(\mathcal{V},\mathcal{A})$ and a vector of weights $\left\{h_{ij}\right\}_{{(i,j)}\in\mathcal{A}}$, the MWRAP consists in finding an arborescence $(\mathcal{V},\mathcal{T})$, with $\mathcal{T}\subset\mathcal{A}$, rooted at a designated node in $G$, that minimizes the sum of weights $\sum_{(i,j)\in\mathcal{T}}h_{ij}$. For more details, we refer the reader to [28, Chapter 6.2]. By definition of the set $\mathbb{S}$, the MWRAP can also be formulated as an integer linear programming problem as the following
@@
Solving (16) as a linear integer problem in every iteration reduces the effectiveness of ADMM as a heuristic approach to our problem. This is attributed to the prolonged convergence time, particularly evident in large-scale networks. In practical scenarios, the MWRAP can be effectively solved by algorithms like Edmond’s algorithm, with a runtime of $O(|\mathcal{A}|+|\mathcal{V}|\log(|\mathcal{V}|))$, as showed in [28, Chapter 6.3].
@@ III-B The proposed algorithm
$s_{k}^{i}$ ::=:= $\left\|\left(\boldsymbol{X},\boldsymbol{b}\right)_{k+1}^{i}-\left(\boldsymbol{X},\boldsymbol{b}\right)_{k}^{i}\right\|,$
$r_{k}^{i}$ ::=:= $\left\|\left(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\lambda}\right)_{k+1}^{i}-\left(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\lambda}\right)_{k}^{i}\right\|.$
@@ IV Numerical experiments
In this section, we demonstrate the efficiency and limitations of our algorithm through various simulations across different scenarios. The section is structured into three subsections. The first subsection examines a simplified network consisting of five nodes, highlighting the resilience of Algorithm 1 under various conditions. We show that our algorithm converges toward the optimal radial solution although the system changes. In the second subsection, Algorithm 1 is implemented on the 33-bus system [4]. We demonstrate the algorithm’s robustness by showcasing its ability to converge despite encountering a line fault during execution. Additionally, we perform a comparative analysis between our algorithm, its centralized counterpart, and the distributed algorithms presented in [17, 18], as well as their enhanced version that will be proposed later. Our findings demonstrate that Algorithm 1 outperforms all others in terms of performance. The final subsection applies Algorithm 1 to a large-scale network operated by SRD, a French transmission system operator.
Unless there is a change, the tolerance parameter $\varepsilon$ for the error $e_{k}$ defined in (44) and the penalization parameter $\delta$ are fixed at $10^{-4}\times|\mathcal{V}|$ and $1$ respectively. The Matlab function quadprog was used to solve problem (26a) Moreover, the MRSAP is solved by implementing Edmonds’ algorithm in Matlab taken from [37].
@@ IV-A Resilience: a toy network
In this subsection, we demonstrate that Algorithm 1 possesses the capability to adapt to dynamic changes within the network environment as they occur in real-time. These adaptations include the detection of modifications such as integrating a new source of energy, switches going offline, or shifts in network traffic patterns. The algorithm subsequently reconfigures the network in response to these changes during its execution without restarting. This process is illustrated in a simple example by considering a graph with a size of $|\mathcal{V}|=5$ and one substation indexed by $1$. The voltage constraint (3) with their corresponding dual updates are dropped. We set $c_{ij}=i+j$, for all $(i,j)\in\mathcal{A}$. Our experience is divided into four phases:
  • from $k=1$ to $k=100$: Algorithm 1 is executed with $\boldsymbol{\rho}^{1}=\boldsymbol{\rho}^{2}=(4,-1,-1,-1,-1)^{T}$. It converges to the optimal solution illustrated in Figure 1a.
  • from $k=100$ to $k=200$: In order to model the failure of the line $(2,1)$ between the $100$th and $200$th iteration, we assign zero values at each iteration to $b_{2,1}^{1}$, $b_{1,2}^{1}$, $b_{1,2}^{2}$, and $b_{2,1}^{2}$ so that this modification is known only to agents $i=1,2$. Algorithm 1 is capable of detecting these line failures and adapting accordingly, ultimately converging to the optimal solution shown in Figure 1b.
  • from $k=200$ to $k=300$: In this phase, we alter the supply-demand vector to $\boldsymbol{\rho}=(4,-1,-2,2,-3)^{T}$ to reflect an increased demand for buses $3$ and $5$, which is compensated by the renewable energy supplied by node $4$. The algorithm seamlessly adapts to these modifications, ultimately converging to the optimal solution within this specific context shown in Figure 1c.
  • from $k=300$ to convergence: During this phase, we assume that the connection between points (1,2) has been restored. Consequently, Algorithm 1 requires modification to ensure it converges towards the optimal solution while adhering to the radiality constraints. These adjustments result in the optimal solution depicted in Figure 1d.
This experience showcases the effectiveness and adaptability of our algorithm in responding to real-time fluctuations, ensuring convergence to a radial solution. The iterative process’s evolution, along with multiple quantities during different phases, is illustrated in Figures 2a-2b.
@@ IV-B Baran and Wu test case
Algorithm 1 has been tested on the Baran and Wu test case [4] which comprises $33$ nodes, $32$ branches, and $5$ tie lines, with a total active and reactive demand equal to $3.715$ MW and $2.24$ MVAr, respectively, and voltage base $12.66$ kV, and a single substation indexed by $1$ linked to a generator, as illustrated in Figure 3. The data of the network can be found in [4]. The optimization process involves comparing the outcome of Algorithm 1 with an initial configuration defined by the inactivation of the following lines: $(25,29)$, $(18,33)$, $(9,15)$, $(8,21)$, $(12,22)$ resulting in a total approximate loss of $183.74$ kW (against $202.7$ kW by using the non-simplified model [9]).
IV-B1 Convergence
As Algorithm 1 is highly influenced by its initial conditions, it underwent ten executions, each with a distinct $\boldsymbol{b}_{0}$. Among these runs, it effectively determined the best solution, activating switches $(8,21)$, $(9,10)$, $(14,15)$, $(28,29)$, and $(32,33)$. This resulted in a loss of $132.28$ kW (against $139.5$ kW by using the non-simplified model [9]), with $\delta$ set at $0.1$. Achieving this outcome required $1349$ iterations, see Figure 4a. Notably, this optimal configuration was attained in only $2$ of the $10$ restarts. In the remaining runs, Algorithm 1 converged but without reaching the optimal solution. See Table I for more details. We chose the $\delta$ value in this specific manner to increase the system’s sensitivity to cost considerations over the quadratic penalization. It is essential to recognize, however, that increasing the value of $\delta$ can potentially sacrifice optimality while simultaneously reducing the number of iterations.
IV-B2 Resilience
We demonstrate the resilience of Algorithm 1 by introducing a fault in the line $(17,18)$ (any other line can be chosen) after the $1200$th iteration. This is done by setting $b_{(17,18)}^{17}=b_{(18,17)}^{17}=b_{(17,18)}^{18}=b_{(18,17)}^{18}=0$, ensuring that the fault is only known locally to agents $17$ and $18$. Under these conditions, Algorithm 1 required $3352$ iterations to converge to the solution with open switches at $(7,8)$, $(10,11)$, $(14,15)$, $(17,18)$, and $(28,29)$, resulting in a loss of $157.8$ kW. The progression of the error $e_{k}$ is illustrated in Figure 4b.

### v2

@@ I Introduction
PDNR methods can be categorized into heuristic [2, 3, 4, 5], meta-heuristic [6, 7, 8] and mathematical optimization-based approaches [9, 10, 11, 12, 13]. Various methods were employed to address the PDNR problem using centralized methods [14]. These methods gather all network data and transmit it back to the control center to determine the reconfiguration solution. However, this approach is vulnerable to single-point failures and cyber-attacks. Furthermore, it can be computationally expensive and unsuitable for large-scale systems.
In response to the limitations of centralized reconfiguration, recent efforts have focused on distributed (or decentralized) algorithms, which offer enhanced scalability, improved fault tolerance, and better adaptability to large-scale systems. In such frameworks, computing agents exchange only limited information with a selected subset of neighboring agents, reducing communication overhead and preserving data privacy. This decentralized structure not only enhances cybersecurity but also lowers the infrastructure costs associated with global coordination. Furthermore, distributed algorithms are inherently robust to individual agent failures, support the integration of renewable energy sources, and enable parallel computation, thereby offering faster convergence and improved scalability compared to centralized approaches [15].
@@
In the context of the distributed approach for the PDNR problem, the multi-agent approach has been explored in several studies, including [16, 17, 18]. However, these approaches are not fully distributed, as they require some or all agents to access global network information, which raises concerns regarding data privacy and introduces vulnerability to single-point failures. Additionally, they often fail to address the need for scalable and efficient solutions suitable for large-scale distribution networks. An alternative method for solving the distributed PDNR problem is the Alternating Direction Method of Multipliers (ADMM), as proposed in [19, 20, 21].
@@
The ADMM algorithm is a primal-dual splitting method initially designed to solve convex optimization problems [22, 23]. ADMM as a general heuristic for solving non-convex problems is mentioned in [22, Ch 9] and this idea is further explored in [24, 25]. Recently, it has garnered attention as a method for discovering approximate solutions for NP-hard problems. Utilizing ADMM for mixed-integer problems offers a significant benefit by requiring only a method to compute the projection onto the discrete constraint set, either exactly or approximately if the computation is costly [24]. While there is no guarantee that ADMM-based methods converge to a global solution for non-convex problems, they are still useful in finding high-quality local solutions.
Regarding the application of ADMM to PDNR problems, several distributed ADMM-based algorithms have been proposed in the literature [19, 20, 21]. In [19], the authors introduced a distributed ADMM algorithm specifically tailored to the PDNR problem. Due to the inherent difficulty of directly handling binary decision variables within the ADMM framework, a relaxation strategy was adopted, as introduced in [24]. In this work, the relaxation approach is presented as a general strategy for handling non-convex constraints within ADMM. However, when applied to PDNR, this method may lead to non-radial network topologies, thus rendering the obtained solution infeasible. The core issue arises in the projection step, which relies on a naive rounding of continuous values to binary decisions (0 or 1) [24], without explicitly enforcing the radiality constraint.
@@
To address this issue, an enhanced ADMM-based algorithm was proposed in [20]. This algorithm incorporates a proximal operator with a penalty parameter based on residual errors to handle binary variables. Nevertheless, there is still no guarantee of achieving a radial solution since the projection step remains on the binary set. In [21], additional conditions were introduced in the projection step to ensure radiality. The authors combined the algorithms from [19] and [20] with their own, selecting the optimal (feasible) solution among the three. Notably, the examination of radiality constraints on binary variables occurs only after the algorithm’s convergence.
Recently, a centralized ADMM-based algorithm was proposed in [13] to solve the PDNR problem. The projection subproblem is addressed exactly using two approaches: the relaxation technique, as used in prior works [19, 20, 21], and a novel variable substitution technique introduced in [13]. Both methods successfully solve the projection step by reformulating it as a Minimum Weight Rooted Arborescence Problem (MWRAP) [26]. However, the substitution technique demonstrated superior performance and favorable properties compared to relaxation. Despite these advances, the proposed centralized algorithm has not yet been extended to the distributed setting.
Contributions: This paper introduces a novel distributed algorithm for solving the PDNR problem, with a particular emphasis on preserving radiality throughout the optimization process. It extends the centralized framework proposed in [13] to a fully decentralized setting, where agents interact solely through local communication with their neighbors. The algorithm is built upon the ADMM framework and leverages a variable substitution technique, initially introduced in [13], which demonstrates superior performance compared to the relaxation-based strategies commonly used in earlier works [19, 20, 21]. Radiality is rigorously enforced at each iteration by reformulating the projection step as a MWRAP with local weights, thereby ensuring the feasibility of the resulting network configuration. Furthermore, we introduce an enhanced version of the relaxation-based algorithms proposed in [19, 20, 21], in which the projection step is solved exactly via a MWRAP, ensuring radiality at each iteration. A comparative analysis on the standard 33-bus distribution test system demonstrates that the proposed variable substitution method consistently outperforms the enhanced relaxation-based approach. This performance advantage, previously established in the centralized setting [13], is now shown to extend to the distributed framework as well.
@@ II The Optimization Problem for the Simplified DistFlow Model
Bold symbols consistently represent vectors or matrices. If $\boldsymbol{x},\boldsymbol{y},\boldsymbol{z}\in\mathbb{R}^{n}$ then $\boldsymbol{z}\in[\boldsymbol{x},\boldsymbol{y}]$ means that $z_{i}\in\left[x_{i},y_{i}\right]$ for all $i\in\{1,\ldots,n\}$. Moreover, $\boldsymbol{x}]_{e}$ denotes the $e$-th element of the vector $\boldsymbol{x}$. The vectors $\boldsymbol{0}_{n}$ and $\boldsymbol{1}_{n}$ denote vectors of size $n$ consisting entirely of zeros and ones, respectively.
@@
$\mathbb{S}=\left\{\begin{array}[]{c}\boldsymbol{b}\in\{0,1\}^{|\mathcal{A}|}:\ \boldsymbol{b}\ \text{forms an arborescence}\\ \text{ rooted at }g\end{array}\right\},$ (9)
@@
$\boldsymbol{A}\boldsymbol{U}]_{e}=U_{i}-U_{j},\quad\text{if }e=(i,j)\in\mathcal{A}.$
Note that a feasible solution $\boldsymbol{b}$ to the above problem defines an arborescence $G_{\boldsymbol{b}}\subseteq G$ (a directed tree) rooted at $g$, whose arcs are the closed switches. The reader can refer to [26, Subsection 2.2] for a detailed definition of an arborescence. Throughout this paper, we assume the feasibility of problem (6), meaning that at least one optimal solution exists.
@@
The voltage drop constraint (3) has been multiplied component-wise with $\boldsymbol{b}$ in (6), making it active solely when $b_{ij}=1$. This is different from the “big-MM” method presented in [10, 29] which includes or excludes these constraints using the inequality:
@@ III-A Review of the MWRAP
Given a directed graph $G=(\mathcal{V},\mathcal{A})$ and a vector of weights $\left\{h_{ij}\right\}_{{(i,j)}\in\mathcal{A}}$, the MWRAP consists in finding an arborescence $(\mathcal{V},\mathcal{T})$, with $\mathcal{T}\subset\mathcal{A}$, rooted at a designated node in $G$, that minimizes the sum of weights $\sum_{(i,j)\in\mathcal{T}}h_{ij}$. For more details, we refer the reader to [26, Chapter 6.2]. By definition of the set $\mathbb{S}$, the MWRAP can also be formulated as an integer linear programming problem as the following
@@
Solving (19) as a linear integer problem in every iteration reduces the effectiveness of ADMM as a heuristic approach to our problem. This is attributed to the prolonged convergence time, particularly evident in large-scale networks. In practical scenarios, the MWRAP can be effectively solved by algorithms like Edmond’s algorithm, with a runtime of $O(|\mathcal{A}|+|\mathcal{V}|\log(|\mathcal{V}|))$, as shown in [26, Chapter 6.3].
@@ III-B The proposed algorithm
$s_{k}^{i}$ $:=$ $\left\|\left(\boldsymbol{X},\boldsymbol{b}\right)_{k+1}^{i}-\left(\boldsymbol{X},\boldsymbol{b}\right)_{k}^{i}\right\|,$
$r_{k}^{i}$ $:=$ $\left\|\left(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\lambda}\right)_{k+1}^{i}-\left(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\lambda}\right)_{k}^{i}\right\|.$
@@ IV Numerical Experiments
In this section, we evaluate the performance and limitations of Algorithm 1 through a series of numerical simulations conducted across various test cases. The objective is to validate the algorithm’s robustness, adaptability, and scalability in both synthetic and realistic scenarios. The section is organized into three parts.
The first subsection investigates a small-scale synthetic network consisting of five nodes, designed to assess the algorithm’s resilience in response to dynamic changes such as topology modifications and variations in supply-demand profiles. We show that Algorithm 1 successfully reconfigures the network and consistently converges to a radial solution under these perturbations.
The second subsection examines the 33-bus distribution test system, originally proposed in [4]. We analyze the convergence behavior of the algorithm under both nominal and faulted conditions. Additionally, a comparative study is presented, contrasting Algorithm 1 with its centralized version proposed in [13] and the distributed algorithms developed in [19, 20]. The comparison focuses on solution quality (optimality), convergence speed, and feasibility.
The third subsection demonstrates the scalability of Algorithm 1 by applying it to a large-scale power distribution network operated by SRD, a French transmission system operator. This case study also includes a fault-handling experiment to assess the algorithm’s real-time resilience.
Unless otherwise specified, the tolerance parameter $\varepsilon$ for the error metric $e_{k}$, defined in (47), is fixed at $10^{-4}\times|\mathcal{V}|$, and the penalization parameter $\delta$ is set to $1$. The quadratic subproblem (29a) is solved using MATLAB’s quadprog function. The MWRAP is solved using an implementation of Edmonds’ algorithm, adapted from [30].
@@ IV-A Resilience: a toy network
IV-A1 Objective
The aim of this subsection is to demonstrate that Algorithm 1 possesses the capability to adapt to dynamic changes within the network environment as they occur in real-time. These adaptations include the detection of modifications such as integrating a new source of energy, switches going offline, or shifts in network traffic patterns. The algorithm subsequently reconfigures the network in response to these changes during its execution without restarting.
IV-A2 Experience
This process is illustrated using a simple example involving a graph of size $|\mathcal{V}|=5$ with one substation indexed by node $1$. The voltage constraint (3) and its corresponding dual updates are omitted. We set the cost coefficients as $c_{ij}=i+j$ for all $(i,j)\in\mathcal{A}$. The experience is divided into four phases:
  • Phase 1: $k=1$ to $k=100$: Algorithm 1 is executed with initial supply-demand vectors $\boldsymbol{\rho}^{i}=(4,-1,-1,-1,-1)^{T}$, $i=1,2$.
  • Phase 2: $k=100$ to $k=200$: A line failure is simulated on edge $(2,1)$ by assigning zero values to $b_{2,1}^{1}$, $b_{1,2}^{1}$, $b_{1,2}^{2}$, and $b_{2,1}^{2}$, with this change locally visible to agents 1 and 2.
  • Phase 3: $k=200$ to $k=300$: The supply-demand vector is modified to $\boldsymbol{\rho}=(4,-1,-2,2,-3)^{T}$, increasing demand on buses 3 and 5 while incorporating renewable generation from node 4.
  • Phase 4: $k=300$ to convergence: The previously failed line $(1,2)$ is restored. The algorithm is then adjusted to accommodate this structural change while enforcing radiality constraints.
IV-A3 Results
After running Algorithm 1, we obtained the following results:
  • In Phase 1, the algorithm converges efficiently to the optimal radial configuration under the initial load distribution, as depicted in Figure 1a.
  • In Phase 2, the algorithm detects the failure of line $(2,1)$ without global knowledge and successfully reconfigures the network. It converges to a new feasible radial structure as shown in Figure 1b.
  • In Phase 3, despite dynamic changes in demand and renewable supply, Algorithm 1 remains robust and converges to the adjusted optimum, illustrated in Figure 1c.
  • In Phase 4, after restoring line $(1,2)$, the algorithm adapts and converges to a new feasible solution that incorporates the restored topology, while preserving radiality. The final configuration is illustrated in Figure 1d.
The evolution of the iterative process, along with several key quantities across different phases, is illustrated in Figures 2a2b. These figures demonstrate that all five agents progressively converge to a common solution, thereby confirming the achievement of consensus.
IV-A4 Conclusion
This example highlights the algorithm’s resilience and adaptability. It efficiently handles real-time changes such as line faults and load shifts while maintaining feasibility, consensus, and radiality throughout all phases.
@@ IV-B Baran and Wu test case
IV-B1 Objective
The objective of this subsection is to evaluate the performance of Algorithm 1 on the 33-bus test system under four distinct scenarios:
  • Scenario 1: Nominal Conditions. This scenario evaluates the algorithm’s ability to reach convergence and consensus when executed under normal conditions, with no structural changes or perturbations throughout the process.
  • Scenario 2: Fault Resilience. A fault is introduced by disabling a specific line in the network. This scenario tests the algorithm’s capacity to detect and adapt to the fault while converging to a new feasible solution that satisfies the network constraints.
  • Scenario 3: Comparison with Centralized Algorithm. The distributed version is compared with the centralized algorithm proposed in [13], with a focus on solution quality, iteration count, and execution time.
  • Scenario 4: Comparison with Other Distributed Methods. The proposed algorithm is benchmarked against the methods developed in [19] and [20], as well as their enhanced versions, which will be introduced in the following subsection.
Algorithm 1 has been tested on the Baran and Wu test case [4] which comprises $33$ nodes, $32$ branches, and $5$ tie lines, with a total active and reactive demand equal to $3.715$ MW and $2.24$ MVAr, respectively, and voltage base $12.66$ kV, and a single substation indexed by $1$ linked to a generator, as illustrated in Figure 3. The data of the network can be found in [4]. The optimization process involves comparing the outcome of Algorithm 1 with an initial configuration defined by the inactivation of the following lines: $(25,29)$, $(18,33)$, $(9,15)$, $(8,21)$, $(12,22)$ resulting in a total approximate loss of $183.74$ kW (against $202.7$ kW by using the non-simplified model [9]). The result for each scenario is presented below.
IV-B2 Convergence
As Algorithm 1 is highly influenced by its initial conditions, it underwent ten executions, each with a distinct $\boldsymbol{b}_{0}$. Among these runs, it effectively determined the best solution, activating switches $(8,21)$, $(9,10)$, $(14,15)$, $(28,29)$, and $(32,33)$. This resulted in a loss of $132.28$ kW (against $139.5$ kW by using the non-simplified model [9]), with $\delta$ set at $0.1$. Achieving this outcome required $1349$ iterations, see Figure 4a. The consensus behavior is also illustrated in Figure 4b.
Notably, this optimal configuration was attained in only $2$ of the $10$ restarts. In the remaining runs, Algorithm 1 converged but without reaching the optimal solution. See Table I for more details. We chose the $\delta$ value in this specific manner to increase the system’s sensitivity to cost considerations over the quadratic penalization. It is essential to recognize, however, that increasing the value of $\delta$ can potentially sacrifice optimality while simultaneously reducing the number of iterations.
IV-B3 Resilience
We demonstrate the resilience of Algorithm 1 by introducing a fault in the line $(17,18)$ (any other line can be chosen) after the $1200$th iteration. This is done by setting $b_{(17,18)}^{17}=b_{(18,17)}^{17}=b_{(17,18)}^{18}=b_{(18,17)}^{18}=0$, ensuring that the fault is only known locally to agents $17$ and $18$. Under these conditions, Algorithm 1 required $2889$ iterations to converge to the solution with open switches at $(7,8)$, $(10,11)$, $(14,15)$, $(17,18)$, and $(28,29)$, resulting in a loss of $157.8$ kW. The progression of the error $e_{k}$ and the agent’s norm behavior are illustrated in Figures 4c and 4d.